OpenStudy (anonymous):

f(x)=2(-x)^2+4(-x)+8

OpenStudy (triciaal):

let y = -x can you factor 2y^2 + 4y + 8

OpenStudy (triciaal):

wait what is the question? are you asked to factorize ?

OpenStudy (anonymous):

well f of x = -x

OpenStudy (anonymous):

2x^2 + 4x + 8 and i plugged in -x but i cant figure it out. im trying to find the value.

OpenStudy (triciaal):

if so i'll proceed divide by 2 to work with smaller numbers and the answer will already be simplified y^2 + 2y + 4 = 0 (y+1)^2= -4 + 1 y + 1 = +/-rt 3 i y = +/-rt 3 i -1 resubstitute y = -x -x =

OpenStudy (triciaal):

well f of x = -x means 2x^2 + 4x + 8 = -x